Fermi National Laboratory


Accelerator Update for November 4, 2005 - November 7, 2005

Friday November 4

The day shift began with Operations monitoring store 4481 and the Recycler stash, with the Tevatron (TeV) conducting an end-of-store study, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, and with NuMI, MiniBooNE, and the Meson MCenter experiments taking beam.

A NuMI expert asked to have their beam turned off at 8:39 AM in preparation for an access later this morning.

Pbar experts began a reverse proton study at 9:05 AM.

A power supply expert reported at 9:33 AM that a MiniBooNE power supply (HV860) was running hotter than it should have been. Power supply technicians began their investigation at 9:37 AM. Operators turned off the MiniBooNE beam for this investigation.

The TeV end-of-store study ended and Operations terminated store 4481 at 9:53 AM.

NuMI experts accessed their experiment to check on their chiller at 10:05 AM.

Power supply experts reported that they had completed their repairs on HV860 at 11:27 AM. MiniBooNE resumed taking beam at 11:29 AM.

Operations reported at 1:16 PM that during the morning RF experts had replaced the Power Amplifier (PA) for a TeV RF station (TRF1) and that the cryo system technicians had replaced a crankshaft and bearings for a TeV sector E3 wet engine and a flywheel for sector B4 wet engine.

At 2:15 PM, a Switchyard (SY) power supply (HP3DS) tripped off on a high temperature indication. Operators contacted an expert.

The reverse protons study ended and Pbar resumed stacking at 2:29 PM.

Power supply experts repaired HP3DS. Operations resumed sending beam to MCenter at 3:30 PM.

Operations put the TeV into shot setup at 4:09 PM.

At 4:23 PM, Operations reported a vacuum cascade in the P1 transfer line. Operators closed vacuum valves and had recovered the area within four minutes.

At 6:09 PM, Operations reported that the Recycler stash had aborted. A power supply tripped off (R:CE411) and took down the rest of the corrector magnet power supplies with it. Unfortunately, this was a very dense stash.

Operations reestablished beam to MCenter at 6:48 PM.

A TeV expert began an injection study at 7:06 PM.

Operations prepared the Recycler at 8:13 PM to stash antiprotons.

Operations put the TeV through a dry squeeze at 9:02 PM.

Operations put the TeV through a wet squeeze at 9:21 PM.

The Recycler stash ended and Pbar resumed stacking. MCenter resumed taking beam a few minutes later.

Operations put the TeV into shot setup at 9:56 PM.

At 11:16 PM, Operations reported that TRF1 tripped off while loading antiprotons into the TeV.

Operations established Recycler only store 4488 at 11:56 PM with an initial luminosity of 65.36E30. Pbar and MCenter resumed normal operations.

Saturday November 5

The midnight shift began with Operations monitoring store 4488, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler began preparing to stash at 2:52 AM. The stash ended around 3:17 AM. Pbar and MCenter resumed normal operations at 3:30 AM.

At 7:01 AM, a Booster RF station (BRF3) tripped off. Operators had the station back on forty minutes later.

The Recycler began preparing to stash at 7:08 AM. The stash ended and Pbar resumed stacking at 7:59 AM. MCenter resumed normal operations at 8:43 AM.

The Recycler began preparing to stash at 3:09 PM. The stash ended. Pbar and MCenter resumed normal operations by 4:11 PM.

Operations terminated store 4488 at 5 PM and put the TeV into shot setup twenty minutes later.

Operations established Recycler only store 4489 at 7:07 PM with an initial luminosity of 121.7E30. Pbar and MCenter resumed normal operations.

At 8:28 PM, the P1 transfer line suffered a vacuum burst.

The Recycler began preparing to stash at 9:04 PM. The stash ended. Pbar and MCenter resumed normal operations by 9:49 PM.

Sunday November 6

The midnight shift began with Operations monitoring store 4489 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler began preparing to stash at 2:25 AM. The stash ended. Pbar and MCenter resumed normal operations by 3:10 AM.

The Recycler began preparing to stash around 7:10 AM. The stash ended. Pbar and MCenter resumed normal operations by 7:53 AM.

CDF reported at 8:38 AM that they had turned off their silicon detector due to bad beam positions shortly after the store was established. Operations contacted a TeV expert and related the experiment’s concerns, but noted that everything looked normal. The expert told Operations to leave things as they were unless they received an alarm.

At 10:31 AM, the NuMI Horn tripped off on a ground fault. Operators informed experts.

The Recycler began preparing to stash at 11:09 AM.

At 11:35 AM, a Booster kicker (MKS02) tripped off on a Klixon indication. Operators investigated.

The stash ended. Pbar and MCenter resumed normal operations around 11:52 AM.

After many resets, operators managed to get MKS02 back on.

At 1:18 PM, Operations reported that the Tevatron Electron Lens (TEL) tripped off. Operators investigated and reset the lens.

NuMI Horn power supply experts had the supply on and running normally by 2 PM.

The Recycler began preparing to stash at 4:05 PM. The stash ended. Pbar and MCenter resumed normal operations around 5:04 PM.

Operations terminated store 4489 at 8:13 PM. Operations put the TeV into shot setup at 8:46 PM.

At 10 PM, a Linac RF station (LRF2) tripped off and wouldn’t reset. Operators called in a Linac expert who had the station back up fifty minutes later.

Operations established store 4490 at 10:24 PM with an initial luminosity of 147.97E30.

The Recycler began preparing to stash at 11:07 PM. The stash ended. Pbar resumed stacking at 11:53 PM.

Monday November 7

The midnight shift began with Operations monitoring store 4490 and the stash, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

MCenter resumed taking beam at 12:47 AM.

The Recycler began preparing to stash at 3:31 AM. The stash ended. Pbar and MCenter resumed normal operations around 4:17 AM.

At 4:41 AM, TRF1 tripped off. Luminosity took a hit, but operators soon stabilized beam positions.

Operations turned off beam to MiniBooNE at 7 AM so that later in the morning experts could access for LCW maintenance.

At 7:50 AM, Operations reported that the TEL tripped off. Operators investigated and reset the supply.

At 8 AM, the P1 transfer line suffered a vacuum burst.

The Plans for Monday

The plans for today are to stack, stash, and store. However, Booster, Main Injector, and Pbar studies will be allowed. The Booster study will disrupt the beam a bit.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/7/2005   email Fermilab